If the interest rate is 6 percent, then the present value of $5,000 received ten years from today is $2,583.34.​

a. True
b. False
Answer #1

The formula used to determine present value is given by :

Present value = Cash flow / (1+r)n

where r is the interest rate and n is the number of years

Thus, Present value = 5000 / (1+6/100) 10

or, Present value = 5000 / (1.06) 10

or, Present value = 5000 / 1.79 = 2793.296

Hence the correct option is false.